OpenStudy (anonymous):

if a rubik's cube has a volume of 384 cubic centimeters, how long is one side of the cube? (recall that the volume of a cube is calculated by L3, where L is the length of one side.

OpenStudy (ranga):

\[\Large L ^{3} = 384\]\[\Large L = \sqrt[3]{384}\]

OpenStudy (mathstudent55):

\(V = L^3 = 384~ cm^3\)

OpenStudy (ranga):

Using your calculator take the cube root of 384